# Multi-modal Machine Learning Analysis of X-ray Absorption Near-Edge Spectra and Pair Distribution Functions: Performance and Interpretability towards Experimental Design

### v1

@@ Abstract
We used off-the-shelf interpretable ML techniques to combine information from multiple heterogeneous spectra: X-ray absorption near-edge spectra (XANES) and atomic pair distribution functions (PDFs), to extract information about local structure and chemistry of transition metal oxides. This approach enabled us to analyze the relative contributions of the different spectra to different prediction tasks. Specifically, we trained random forest models on XANES, PDF, and both of them combined, to extract charge (oxidation) state, coordination number, and mean nearest-neighbor bond length of transition metal cations in oxides. We find that XANES-only models tend to outperform the PDF-only models for all the tasks, and information from XANES often dominated when the two inputs were combined. This was even true for structural tasks where we might expect PDF to dominate. However, the performance gap closes when we used species-specific differential PDFs (dPDFs) as the inputs instead of total PDFs. Our results highlight that XANES contains rich structural information and may be further developed as a structural probe. Our interpretable, multimodal approach is quick and easy to implement when suitable structural and spectroscopic databases are available. This approach provides valuable insights into the relative strengths of different modalities for a practical scientific goal, guiding researchers in their experiment design tasks such as deciding when it is useful to combine complementary techniques in a scientific investigation.
@@ 1 Introduction
Extracting local atomic structure is crucial for characterizing materials and understanding their properties [egami;b;utbp12]. This is often achieved by analyzing scattering or spectroscopy data, but sometimes it is not possible to obtain the relevant details from one method alone due to limited data quality or problem complexity [billi;s07, billi;p10]. Researchers often fit an internal model of a system to experimental observations via optimization [mcgreevy1988reverse, billingeRealspaceRietveldFull1998b]. For complex cases, it may be desirable to combining information from two or more different experiments to obtain the desired information. This is common in scientific literature where many characterizations are done on the same sample, but a greater challenge is to use multiple experiments to refine parameters from a single model, a process known as complex modeling [juhas;aca15] or multimodal data fusion [lahat2015multimodal]. This has proved more challenging than expected due to the heterogeneous nature of information in different experiments and possible incompatibility of systematic errors [farrowunpublished]. Few attempts have been made to combine data from two different structural probes in a rational co-refinement. One approach combined PDF and small-angle x-ray scattering (SAXS) [farrow2014robust], while another merged extended X-ray absorption fine structure (EXAFS) and total scattering/PDF data in reverse Monte Carlo (RMC) refinements [tucke;jpcm07, krayzman2008simultaneous]. However in general, traditional methods still struggle to integrate heterogeneous information from different experiments, in large part because we lack a priori knowledge on how to weight the contributions from each measurement in the cost function.
@@
Both XANES and PDF are powerful and popular techniques to probe local atomic environments [yano2009x, egami;b;utbp12]. They are also both well-suited to a database approach. Theoretical X-ray absorption spectra are relatively inexpensive to obtain using FEFF [rehr2010parameter] and are a part of the Materials Project (MP) database [mathew2018high, jain2013commentary], and PDFs are straightforward to compute with Diffpy-CMI [juhas;aca15, billinge2023atomic] from the atomic coordinates. These two methods present an interesting pair of modalities because their information content may overlap despite their differing physical origins. X-ray absorption at an element’s absorption edge produces XANES features that reflect the electronic structure of the resonant atom, while X-ray diffraction and scattering generate PDF patterns that reveal local geometrical arrangements of atoms. Understanding the information content in XANES and PDF, and their complementarity, can inform experimental design: when to use which technique and when it is useful to combine them. Experiments, particularly conducted at beamlines at national user facilities, are valuable resources. Gaining early insights how to design experimental campaigns for the optimal information yield can help the scientific community make best use of expensive resources. We show how interpretable ML can do this straightforwardly by revealing information that off-the-shelf nonlinear models can extract from different spectrum types.
@@
One major advantage of using ML is the ability to extract obscure information from complex spectral data. Many ML models have been developed and implemented for XANES analysis. Review articles by Timoshenko & Frenkel [timoshenko2019inverting] and Penfold et al. [penfold2024machinelearning] and references therein cover recent applications of ML for prediction and for understanding structure-spectra relationship from XANES. Supervised ML techniques such as convolutional neural networks (CNNs) and random forest models were proven effective at determining the absorbing atom’s environment including coordination number [timoshenko2017supervised, carbone_classification_2019], oxidation state [zheng_random_2020], and nearest-neighbor bond lengths [torrisi_random_2020]. More recently, with generative AI on the rise, diffusion models were employed to generate 3D structure candidates of amorphous carbon conditioned on XANES [kwon2023spectroscopyguided].
@@
With emphasis on interpretability of the results, some of these studies identified regions of the XANES spectra that contributed the most insights for the model prediction task, either by including only the pre-edge region as input [carbone_classification_2019] or by evaluating the feature importance of random forest models [torrisi_random_2020, zheng_random_2020]. Tetef et al. did similar analyses to compare information content in XANES and valence-to-core x-ray emission spectra (VtC-XES) [tetef2021unsupervised, tetef2022informed]. Alternatively, Guda et al. featurized XANES based on spectral descriptors, such as edge position and energy, and analyzed their correlations with properties [guda_understanding_2021]. Additionally, unsupervised ML techniques, such as principal component analysis (PCA), have been employed to examine the structure-spectrum relationship within a low-dimensional latent space [guda_understanding_2021, tetef2021unsupervised, liang2023decoding].
@@
Compared to XANES, fewer ML models have been developed to extract structural information from PDF. Although determining structure directly from PDF (the nanostructure inverse problem [juhas;aca15]) remains a challenging task, ML has enabled progress on crystals and nanoparticles. For crystal structures, Liu et al. developed a CNN for space group identification [liu2019using]. Trained on calculated PDFs, their model was robust [lan2022robustness] against different parameter choices in PDF calculations [billinge2023atomic]. For nanoparticles, Kjær and colleagues developed DeepStruc, a graph-based conditional variational autoencoder (CVAE) algorithm [kjaerDeepStrucStructureSolution2023] that can solve monometallic nanoparticle structures from calculated and experimental PDFs. Other ML prediction tasks from PDF include that of Ref. [zhang2023pair], where unsupervised techniques including PCA and non-negative matrix factorization (NMF) were applied to feature engineer PDF for prediction of defect concentrations in Ti oxides. On the other hand, some studies have integrated ML methods to improve the model refinements, the conventional approach for PDF analysis, by using gradient boosted decision trees [anker2022extracting] to select the best starting model for fitting, or combining ML and DFT [klove2023machinelearningbased] to automate the model selection and improve the optimization algorithm. NMF has been shown effective for decomposition of PDFs into components that resemble partial (differential) PDFs of different chemical components of the system [liu2021validation] and can be easily implemented [thatcher2022nmfmapping]. More recently, with generative AI, Guo et al. developed models to solve a structure solution from x-ray diffraction (XRD) data from nanoparticles, outputting a continuous electron density map [guo2023endtoend] or predicting a unit cell with discrete atomic coordinates [guoinitio].
@@
Unlike traditional methods, ML can readily combine heterogeneous inputs [baltruvsaitis2018multimodal, hastie2009elements], made even easier when using techniques like random forests [breiman2001random] that can flexibly handle diverse input types off-the-shelf with minimal numerical issues. Very recently, applications of ML in materials science started incorporating multimodal inputs to improve predictive performance. On XANES, Carbone et al. showed that combining XANES spectra from multiple species (C, H, and N) improved the model’s performance in predicting chemical motifs of small organic molecules [carbone2024accurate]. Szymanski et al. combined XRD patterns and PDFs as input into their neural networks and improved the accuracy for space group identification [szymanski_integrated_2024-2]. They found that models trained on PDF or XRD separately often failed on different samples, but combining them enhanced performance. We note in passing that combining inputs has been done in the context of conventional modeling, for example the RMCprofile method [tucke;jpcm07] where combined PDF and XRD was needed to obtain models that were sufficiently well ordered. In this case the inputs are two representations of the same information related through a Fourier transform, albeit where the information is differently weighted in the two representations. However, combining inputs with different physical origins and properties is more challenging, limiting the number of such studies. Unlike prior multimodal ML models [carbone2024accurate, szymanski_integrated_2024-2] that used complementary inputs with similar origins and did not explore interpretability, our work aims to overcome heterogeneity between any arbitrary modalities and investigate how information is balanced between them, hence we focus on interpretability and relative importance of the input features.
@@
In this work, we investigate with interpretable ML the information content and complementarity of XANES and PDF. We also explore whether a multimodal approach results in improvement, i.e., whether using XANES and PDF together as inputs to the model improves predictions of targets that describe local materials structure. The system we chose for the demonstration was transition metal (Ti, Mn, Fe, Cu) oxides. We curated four separate datasets, corresponding to the four metals, consisting of structures, theoretically calculated XANES spectra [mathew2018high] (K-edge of the transition metal atom) available on the Materials Project (MP) database [jain2013commentary], and PDFs computed with Diffpy-CMI [juhas;aca15, billinge2023atomic]. Separately, on each dataset we trained random forest models [breiman2001random] to extract the metal’s charge (oxidation) state, coordination number, and average nearest-neighbor bond lengths. These targets encode properties of the local atomic environment of the transition metal atoms, ranging from electronic to purely structural characteristics. Similar sets of prediction tasks have been performed on XANES and used to evaluate contributions of different features and regions of the spectra, particularly by Torrisi et al. [torrisi_random_2020] and Guda et al. [guda_understanding_2021], but these tasks have not been the focus of previous PDF studies with ML. This work offers not only the database-level analysis of PDFs but also a systematic comparison of XANES and PDFs across various prediction tasks, providing valuable insights for practitioners of both experimental techniques.
@@
The information content, or fitness to a particular problem, depends on both the input and the output of interest. We use random forest models because off-the-shelf implementations allow us to study the feature importance during training, giving us an unbiased assessment of which modality and where in the XANES and PDF spectra contribute the most to reducing training loss for each prediction task. When combining XANES and PDF, feature importance plots also reveal how the information is balanced between the two data sources and when one input dominates.
@@
Here our main result is to compare the random forest model performances when trained on the different inputs: XANES, PDF, and both of them combined, on the three prediction tasks: charge (oxidation) state, coordination number, and mean nearest-neighbor bond length. Surprisingly, models trained on XANES outperformed those trained on PDFs in most cases, even for structural targets like bond length. We showed that this advantage went away when site-specific differential-PDFs (dPDFs) were used as the PDF inputs, indicating that the advantage of XANES comes from its species sensitivity (assuming that the edge energies are well-separated from other species that are present, which is commonly the case for K-edges of transition metals). In 3/4 cases, combining XANES and dPDFs improved the model performances over using XANES alone, and in one case outperformed dPDF. This suggests that dPDFs contribute complementary local structural information to the model.
@@
This study illustrates the power of ML models to easily combine heterogeneous measured spectra that contain distinct and complementary information to accomplish a classification or regression task. In multi-modal analysis, one of the greatest challenges is to know how to weight the contributions of the different inputs. Our approach, of investigating the feature importance of the models in their tasks, may suggest how practitioners may approach the weighting of two modalities. These insights enhance our understanding of the data and enable us to assess the relative importance of different measurements for success in a particular task. Our investigation into the differing performance of XANES, PDF, and dPDF provides a potential template for future studies which seek to understand the difference in utility of different spectral modes.
@@ 2 Results
2.1 Charge state classification results
First we report results for the classification of charge (oxidation) state from XANES and PDF separately. A random forest model was trained to classify three charge states, $CS=+2$, +3, and +4 for Ti, Mn, and Fe, and $CS=+1$, +2, and +3 for Cu. Due to class imbalances in the charge state, we use the F1 score to compute performance metrics as the standard accuracy metric can be misleading. We also use a weighted mean that takes into account the class proportion in the dataset. The procedure for data procurement, labeling, and model training is described in detail in the Methods section, Sec. 4.
Figure 1: Charge state classification results: (a) test F1 scores for XANES, PDF, and XANES+PDF models, respectively, from left to right on the four datasets: Ti, Mn, Fe, and Cu. Black horizontal lines are F1 scores obtained by a trivial classifier that labels all samples as the majority class. Error bars on the test scores represent one standard deviation. Feature importance plots are shown for (b) Ti and (c) Fe datasets, for (i) XANES features, (ii) PDF features, and (iii) XANES+PDF features. Dotted lines mark the locations of prominent features of XANES- and PDF-only feature importances. To better understand if the important features change in the models trained on XANES+PDF, the same lines are drawn on the combined XANES+PDF importance plot (iii) to highlight the changes.
As Fig. 1(a) shows, XANES and PDF models performed well giving F1 scores between 0.83 and 0.96, and 0.76 and 0.86, respectively, depending on the transition metal. These scores are well above the baseline scores computed for a trivial classifier that labels all samples as the majority class, shown in the figure as black horizontal lines. These baselines scores are what we might expect if the input data did not contain any predictive information for the task. The baseline scores are high for Ti (0.73) and Cu (0.63) because these datasets are heavily imbalanced, hence labeling all samples as the dominant class yields high scores, but these baseline scores are still significantly lower than the F1 scores achieved by our XANES and PDF models. This indicates that both PDF and XANES data contain rich information that is predictive of the charge state of the ions, as expected.

### v2

# Interpretable Multimodal Machine Learning Analysis of X-ray Absorption Near-Edge Spectra and Pair Distribution Functions

@@ Abstract
We used interpretable machine learning to combine information from multiple heterogeneous spectra: X-ray absorption near-edge spectra (XANES) and atomic pair distribution functions (PDFs) to extract local structural and chemical environments of transition metal cations in oxides. Random forest models were trained on simulated XANES, PDF, and both combined to extract oxidation state, coordination number, and mean nearest-neighbor bond length. XANES-only models generally outperformed PDF-only models, even for structural tasks, although using the metal’s differential PDFs (dPDFs) instead of total PDFs narrowed this gap. When combined with PDFs, information from XANES often dominates the prediction. Our results demonstrate that XANES contains rich structural information and highlight the utility of species-specificity. This interpretable, multimodal approach is quick to implement with suitable databases and offers valuable insights into the relative strengths of different modalities, guiding researchers in experiment design and identifying when combining complementary techniques adds meaningful information to a scientific investigation.
@@ 1 Introduction
Extracting local atomic structure is crucial for characterizing materials and understanding their properties [1]. This is often achieved by analyzing scattering or spectroscopy data, but sometimes it is not possible to obtain the relevant details from one method alone due to limited data quality or problem complexity [2, 3]. Researchers often fit an internal model of a system to experimental observations via optimization [4, 5]. For complex cases, it may be desirable to combining information from two or more different experiments to obtain the desired information. This is common in scientific literature where many characterizations are done on the same sample, but a greater challenge is to use multiple experiments to refine parameters from a single model, a process known as complex modeling [6] or multimodal data fusion [7]. This has proved more challenging than expected due to the heterogeneous nature of information in different experiments and possible incompatibility of systematic errors [8]. Few attempts have been made to combine data from two different structural probes in a rational co-refinement. One approach combined pair distribution functions (PDFs) [1, 9] and small-angle x-ray scattering (SAXS) [10], while another merged extended X-ray absorption fine structure (EXAFS) and total scattering/PDF data in reverse Monte Carlo (RMC) refinements [11, 12]. However in general, traditional methods still struggle to integrate heterogeneous information from different experiments, in large part because we lack a priori knowledge on how to weight the contributions from each measurement in the cost function.
@@
Both XANES and PDF are powerful and popular techniques to probe local atomic environments [13, 1]. They are also both well-suited to a database approach. Theoretical X-ray absorption spectra are relatively inexpensive to obtain using FEFF [14] and are a part of the Materials Project (MP) database [15, 16], and PDFs are straightforward to compute with Diffpy-CMI [6, 9] from the atomic coordinates. These two methods present an interesting pair of modalities because their information content may overlap despite their differing physical origins. X-ray absorption at an element’s absorption edge produces XANES features that reflect the electronic structure of the resonant atom, while X-ray diffraction and scattering generate PDF patterns that reveal local geometrical arrangements of atoms. Understanding the information content in XANES and PDF, and their complementarity, can inform experimental design: when to use which technique and when it is useful to combine them. Experiments, particularly conducted at beamlines at national user facilities, are valuable resources. Gaining early insights how to design experimental campaigns for the optimal information yield can help the scientific community make best use of expensive resources. We show how interpretable ML can do this straightforwardly by revealing information that off-the-shelf nonlinear models can extract from different spectrum types.
@@
One major advantage of using ML is the ability to extract obscure information from complex spectral data. Many ML models have been developed and implemented for XANES analysis. Review articles by Timoshenko & Frenkel [17] and Penfold et al. [18] and references therein cover recent applications of ML for prediction and for understanding structure-spectra relationship from XANES. ML techniques that rely on labeled example data (supervised learning) have been applied to inverse problems with XANES, such as mapping from spectral data to material properties. Such models include but are not limited to heavily parameterized function approximators like convolutional neural networks (CNNs) and decision tree ensembles (i.e. random forest models) [19, 20]. Supervised ML techniques such as convolutional neural networks (CNNs) and random forest models were proven effective at determining the absorbing atom’s environment including coordination number [21, 22, 23], oxidation state [24], and nearest-neighbor bond lengths [24]. More recently, with generative AI on the rise, diffusion models were employed to generate 3D structure candidates of amorphous carbon conditioned on XANES [25].
@@
With emphasis on interpretability of the results, some of these studies identified regions of the XANES spectra that contributed the most insights for the model prediction task, either by including only the pre-edge region as input [22] or by evaluating the feature importance of random forest models [24, 23]. Tetef et al. did similar analyses to compare information content in XANES and valence-to-core x-ray emission spectra (VtC-XES) [26, 27]. Alternatively, Guda et al. featurized XANES based on spectral descriptors, such as edge position and energy, and analyzed their correlations with properties [28]. Additionally, unsupervised ML techniques, such as principal component analysis (PCA), have been employed to examine the structure-spectrum relationship within a low-dimensional latent space [28, 26, 29].
@@
Compared to XANES, fewer ML models have been developed to extract structural information from PDF. Although determining structure directly from PDF (the nanostructure inverse problem [6]) remains a challenging task, ML has enabled progress on crystals and nanoparticles. For crystal structures, Liu et al. developed a CNN for space group identification [30]. Trained on calculated PDFs, their model was robust [31] against different parameter choices in PDF calculations [9]. For nanoparticles, Kjær and colleagues developed DeepStruc, a graph-based conditional variational autoencoder (CVAE) algorithm [32] that can solve monometallic nanoparticle structures from calculated and experimental PDFs. Other ML prediction tasks from PDF include that of Ref. [33], where unsupervised techniques including PCA and non-negative matrix factorization (NMF) were applied to feature engineer PDF for prediction of defect concentrations in Ti oxides. On the other hand, some studies have integrated ML methods to improve the model refinements, the conventional approach for PDF analysis, by using gradient boosted decision trees [34] to select the best starting model for fitting, or combining ML and DFT [35] to automate the model selection and improve the optimization algorithm. NMF has been shown effective for decomposition of PDFs into components that resemble partial (differential) PDFs of different chemical components of the system [36] and can be easily implemented [37]. More recently, with generative AI, Guo et al. developed models to solve a structure solution from x-ray diffraction (XRD) data from nanoparticles, outputting a continuous electron density map [38] or predicting a unit cell with discrete atomic coordinates [39].
@@
Unlike traditional methods, ML can readily combine heterogeneous inputs [40, 41], made even easier when using techniques like random forests [20] that can flexibly handle diverse input types off-the-shelf with minimal numerical issues. Very recently, applications of ML in materials science started incorporating multimodal inputs to improve predictive performance. On XANES, Carbone et al. showed that combining XANES spectra from multiple species (C, H, and N) improved the model’s performance in predicting chemical motifs of small organic molecules [42]. Szymanski et al. combined XRD patterns and PDFs as input into their neural networks and improved the accuracy for space group identification [43]. They found that models trained on PDF or XRD separately often failed on different samples, but combining them enhanced performance. We note in passing that combining inputs has been done in the context of conventional modeling, for example the RMCprofile method [11] where combined PDF and XRD was needed to obtain models that were sufficiently well ordered. In this case the inputs are two representations of the same information related through a Fourier transform, albeit where the information is differently weighted in the two representations. However, combining inputs with different physical origins and properties is more challenging, limiting the number of such studies. Unlike prior multimodal ML models [42, 43] that used complementary inputs with similar origins and did not explore interpretability, our work aims to overcome heterogeneity between any arbitrary modalities and investigate how information is balanced between them, hence we focus on interpretability and relative importance of the input features.
@@
In this work, we investigate with interpretable ML the information content and complementarity of XANES and PDF. We also explore whether a multimodal approach results in improvement, i.e., whether using XANES and PDF together as inputs to the model improves predictions of targets that describe local materials structure. The system we chose for the demonstration was transition metal oxides. We curated four separate datasets, corresponding to the four metals (Ti, Mn, Fe, Cu), consisting of structures, theoretically calculated XANES spectra [15] (K-edge of the transition metal atom) available on the Materials Project (MP) database [16], and PDFs computed with Diffpy-CMI [6, 9]. Separately, on each dataset we trained random forest models [20] to extract the metal’s oxidation state, coordination number, and average nearest-neighbor bond lengths. These targets encode properties of the local atomic environment of the transition metal atoms, ranging from electronic to purely structural characteristics. Similar sets of prediction tasks have been performed on XANES and used to evaluate contributions of different features and regions of the spectra, particularly by Torrisi et al. [24] and Guda et al. [28], but these tasks have not been the focus of previous PDF studies with ML. This work offers not only the database-level analysis of PDFs but also a systematic comparison of XANES and PDFs across various prediction tasks, providing valuable insights for practitioners of both experimental techniques.
@@
The information content, or fitness to a particular problem, depends on both the input and the output of interest. We chose to use random forest models because off-the-shelf implementations allow us to study the feature importance [20] during training. Feature importance is a score assigned to each input feature based on how much it contributes to the trained model’s prediction, giving us an unbiased assessment of which modality and where in the XANES and PDF spectra contribute the most to reducing training loss for each prediction task. When combining XANES and PDF, feature importance plots also reveal how the information is balanced between the two data sources and when one input dominates.
@@
Here our main result is to compare the random forest model performances when trained on the different inputs: XANES, PDF, and both of them combined, on the three prediction tasks: oxidation state, coordination number, and mean nearest-neighbor bond length. We used F1 scores (harmonic mean of precision and recall) as a metric for classification tasks (oxidation state and coordination number) and root mean square errors (RMSEs) for the bond length regression. Surprisingly, models trained on XANES outperformed those trained on PDFs in most cases, even for structural targets like bond length. We showed that this advantage went away when site-specific differential-PDFs (dPDFs) were used as the PDF inputs, indicating that the advantage of XANES comes from its species sensitivity (assuming that the edge energies are well-separated from other species that are present, which is commonly the case for K-edges of transition metals). In 3/4 cases, combining XANES and dPDFs improved the model performances over using XANES alone, and in one case outperformed dPDF. This suggests that dPDFs contribute complementary local structural information to the model.
@@
This study illustrates the power of ML models to easily combine heterogeneous measured spectra that contain distinct and complementary information to accomplish a classification or regression task. In multimodal analysis, one of the greatest challenges is to know how to weight the contributions of the different inputs. Our approach, of investigating the feature importance of the models in their tasks, may suggest how practitioners may approach the weighting of two modalities. These insights enhance our understanding of the data and enable us to assess the relative importance of different measurements for success in a particular task. Our investigation into the differing performance of XANES, PDF, and dPDF provides a potential template for future studies which seek to understand the difference in utility of different spectral modes.
@@ 2 Results
2.1 Oxidation state classification
First we report results for the classification of oxidation state from XANES and PDF separately. A random forest model was trained to classify three oxidation states, $CS=+2$, +3, and +4 for Ti, Mn, and Fe, and $CS=+1$, +2, and +3 for Cu. Due to heavy class imbalances in the oxidation state, we use the F1 score as a performance metric as the standard accuracy score (true positive rate) alone can be misleading. The F1 score of each class is the harmonic mean of the precision $P$ and the recall $R$, ranging from 0 (worst) to 1 (best), where
$P$ $=\frac{TP}{TP+FP}$ (1)
$R$ $=\frac{TP}{TP+FN}$ (2)
$F1$ $=\frac{2PR}{P+R}=\frac{2TP}{2TP+FP+FN},$ (3)
and $TP$, $FP$, and $FN$ are true positive, false positive, and false negative rates, respectively. We compute F1 scores for each class and a weighted mean across all classes that takes into account the proportion of the class in the dataset. The test scores shown in all figures are the weighted mean F1 scores. As a basis for comparison and evaluation of our models, we also compute baseline scores (weighted mean F1) achieved by a trivial classifier that, naively, always predicts the modal class. A well-performing model should achieve a significantly higher F1 score than this baseline. The procedure for data procurement, labeling, and model training is described in detail in the Methods section, Sec. 4.
Figure 1: Oxidation state classification results: (a) test F1 scores for XANES (left bar of each triplet, no pattern), PDF (middle bar with circular pattern), and XANES+PDF models (right bar with striped pattern). Results are shown for the four datasets, from left to right: Ti (orange), Mn (green), Fe (blue), and Cu (magenta). Black horizontal lines indicate F1 scores from a trivial classifier that labels all samples as the modal class. Error bars on the test scores represent one standard deviation. Feature importance plots are shown for (b) Ti and (c) Fe datasets, for (i) XANES features, (ii) PDF features, and (iii) XANES+PDF features. In (i) and (ii), the XANES and PDF spectra are plotted in light gray, with scales defined by the secondary $y-$axis on the right. The dark gray line represents the average spectrum. Vertical dotted lines mark the locations of prominent features from the XANES- and PDF-only importance plots. To highlight changes in feature importance for the combined XANES+PDF models, the same lines are shown in (iii) for comparison.
As Fig. 1(a) shows, XANES and PDF models performed well giving F1 scores between 0.83 and 0.96, and 0.76 and 0.86, respectively, depending on the transition metal. These scores are well above the baseline F1 scores computed for a trivial classifier that labels all samples as the modal class, shown in the figure as black horizontal lines. These baselines scores are what we might expect if the input data did not contain any predictive information for the task. The baseline scores are high for Ti (0.73) and Cu (0.63) because these datasets are heavily imbalanced, hence labeling all samples as the dominant class yields high scores, but these baseline scores are still significantly lower than the F1 scores achieved by our XANES and PDF models. This indicates that both PDF and XANES data contain rich information that is predictive of the oxidation state of the ions, as expected.